OpenStudy (anonymous):

Determine the factors of x2 − 6x − 8. Prime (x + 4)(x − 2) (x − 4)(x + 2) (x + 8)(x − 1)

OpenStudy (mertsj):

What are the factors of -8 whose sum is -6?

OpenStudy (mertsj):

-8=-1x 8, -1+8=-7 That one does not work. -8=1 x -8, 1+-8=7 That does not work either. -8=-2 x 4, -2+4=2 That does not work. -8=2 x -4, 2+ -4=-2 That does not work either. There are no other factors of -8 so the given trinomial is prime.

OpenStudy (anonymous):

-1+8=-7

OpenStudy (mertsj):

Thank you for pointing out that typographical error.

OpenStudy (anonymous):

np